A stationary hockey puck with a mass of 160 grams has been hit with a hockey stick that applies a force of 19 N.  If the puck is in contact with the hockey stick for a distance of 0.18 m, how much momentum has been given to the puck as a result of the hit?


1
Expert's answer
2021-05-05T10:53:45-0400

"F=ma\to a=F/m=19/0.16=118.75\ (m/s^2)"


"s=v^2/(2a)\to v=\sqrt{2as}=\sqrt{2\cdot118.75\cdot0.18}=6.54\ (m/s)"


"p=mv=0.16\cdot6.54=1.05\ (kg\cdot m/s)" . Answer